Instructions Traxonia Railroad Inc. has three regional divisions organized as profit centers. The chief executive officer (CEO) evaluates divisional performance, using income from operations as a percent of revenues. The following quarterly inoome and expense accounts were provided from the trial balance as of December 31, 2016: Revenues-East $870,000 Revenues-West 1,034,000 Revenues-Central 1,880,000 Operating Expenses-East 565,700 Operating Expenses-West 621,360 Operating Expenses-Central 1,174,660 Corporate Expenses-Shareholder Relations 150,000 Corporate Expenses-Customer Support 350,000 Corporate Expenses-Legal 264:000 General Corporate Officers' Salaries 281,000 The company operates three service departments: Shareholder Relations, Customer Support, and Legal. The Shareholder Relations Department oonducts a variety of services for shareholders of the company. The Customer Support Department is the company's point of contact for new service,, complaints, and requests for repair. The department believes that the number of customer contacts is an activity base for this work. The Legal Department provides legal services for division management. The department beleves that the number of hours billed is an activity base for this work. The following additional information has been gathered: East West Central Number of customer contacts 3,000 4,000 7,000 Number of hours billed 1,300 2,160 2,040
Profit center responsibility reporting for a service company Instructions Quarterly Income Statements Final Questions Instructions Quarterly Income Statements Tra For the Quarter Ended December 31, 2016 aр 1 East West Central 2 Revenues $870,000.00 $1,034,000.00 $1,880,000.00 Operating expenses 565,700.00 621,360.00 1,174,660.00 3 4 Income from operations before service department charges 5Less service department charges: Customer Support Legal Total service department charges Income from operations The company operates three service departments: Shareholder Relations, Customer Support, and Legal. The Shareholder Relations Department conducts a varie shareholders of the company. The Customer Support Department is the company's point of contact for new service, complaints, and requests for repair. The depar that the number of customer contacts is an activity base for this work. The Legal Department provides legal services for division management. The department beli number of hours billed is an activity base for this work. The following additional information has been gathered: East West Central
2 Compute the proft margin for each division Note: Enter percentage rounded two decimal places (eg 22547 is 22.50%) Division Profit Margin East Division West Division Central Division Now identity the most successful division acoording to the proft margin 3 What would you include in a recommendation to the CEO for a better method for evaluating the performance of the divisions? What is a major weakness of the present method? A major weakness of the present method is that O there is no weakness. The present method works well a ful year's income is needed for assessment the service department charges are incomectly alocated. the assets invested in each division are not considered. Ononfinancial drivers are not identfied
Instructions Quarterly Income Statements Final Questions there is no weakness. The present method works well. a full year's income is needed for assessment. the service department charges are incorrectly allocated. the assets invested in each division are not considered. nonfinancial drivers are not identified. Which of the following methods would better evaluate divisional performance? Check all that apply None of these. The present method works well Considering residual income (income frorkoperations less a minimal return on divisional assets) Computing the rate of return on investment (income from operations divided by divisional assets) Completing a balanced scorecard for each service department Utilizing transfer pricing between divisions Including direct and indirect operating expenses for each division Focusing on controllable revenues and expenses
Answer #1

1.

1 East West Central
2 Revenues 870000 1034000 1880000
3 Operating expenses 565700 621360 1174660
4 Income from operations before service department charges 304300 412640 705340
5 Less service department charges:
6 Customer support 75000 100000 175000
7 Legal 62400 103680 97920
8 Total service department charges 137400 203680 272920
9 Income from operations 166900 208960 432420

Working:

Customer support:
East: $350000 x 3000/(3000 + 4000 + 7000) = $350000 x 3000/14000 = $75000
West: $350000 x 4000/14000 = $100000
Central: $350000 x 7000/14000 = $175000
Legal:
East: $264000 x 1300/(1300 + 2160 + 2040) = $264000 x 1300/5500 = $62400
West: $264000 x 2160/5500 = $103680
Central: $264000 x 2040/5500 = $97920

2.

Division Profit Margin
East Division 19.18%
West Division 20.21%
Central Division 23.00%

Most successful division according to profit margin: Central

Working:

Profit margin = Income from operations/Revenues

East: $166900/$870000 = 19.18%

West = $208960/$1034000 = 20.21%

Central: $432420/$1880000 = 23.00%

3. The assets invested in each division are not considered.

4. Considering residual income (income from operations less a minimum return on divisional assets)

Computing the rate of return on investment (income from operations divided by divisional assets)